OpenStudy (anonymous):

derivative y=In^4(8x-x^3)

OpenStudy (anonymous):

what grade r u in?

OpenStudy (anonymous):

first

OpenStudy (anonymous):

-_- we mean

OpenStudy (anonymous):

can u give me a medal

OpenStudy (callisto):

@Suzan987654321 at least you should help aga solve the problem before you ask for medal?

OpenStudy (anonymous):

u right callisto! :)

OpenStudy (callisto):

is the expression \[y = \ln^{4}(8x-x^3)\]?

OpenStudy (anonymous):

is this a final answer?

OpenStudy (callisto):

no i'm asking is that the expression of the question

OpenStudy (anonymous):

oohh sorry,yest thats the one

OpenStudy (callisto):

\[dy/dx =d/dx [\ln^{4}(8x-x^3)]\]\[ =d/dx [\ln(8x-x^3)]^4\]\[ = 4[ln(8x-x^3)]^3[1/(8x-x^3)](8-2x^2) \]\[ = 4(8-2x^2)[ln(8x-x^3)]^3/(8x-x^3)\]\[ = 8(4-x^2)[ln(8x-x^3)]^3/(8x-x^3)\]

OpenStudy (anonymous):

thanx,can u please explane how did u get this one 4[ln(8x−x3)]3[1/(8x−x3)](8−2x2)